# Degenerate parabolic equations in divergence form: fundamental solution and Gaussian bounds

### v2

@@ 4.3. Conservation property
When $R\to\infty$, the term in the left hand side tends to $\int_{\mathbb{R}^{n}}\overline{\psi(x)}\ \mathrm{d}\omega(x)$ and the first term in the right hand side tends to $\int_{\mathbb{R}^{n}}\overline{(\tilde{\Gamma}(s,t)\psi)(x)}\ \mathrm{d}\omega(x)$ by Corollary 4.2. It remains to show that the second term in the right hand side tends to 00 as $R\to\infty$. For $R>1$, we have
@@
5. Gaussian upper bounds and Moser’s $L^{2}$-$L^{\infty}$ estimates on local weak solutions
In this section, we prove the equivalence between Gaussian upper bounds for $\mathcal{H}$ and Moser’s $L^{2}$-$L^{\infty}$ estimates for $\mathcal{H}$ and $\mathcal{H}^{\star}$. We will first define these notions rigorously.
@@ 5.1. Gaussian upper bounds and the generalized fundamental solution
See [CUR14, Rem. 3], where it is proven that there is a constant $C=C(D,k_{0})>0$ verifying
@@
5.2. Moser’s $L^{2}$-$L^{\infty}$ estimates
@@
We say that $\mathcal{H}$, respectively $\mathcal{H}^{\star}$, satisfies Moser’s $L^{2}$-$L^{\infty}$ estimates if there exists a constant $B>0$ such that that for all $R>0$, $(t_{0},x_{0})\in\mathbb{R}^{1+n}$, and all local weak solutions of $\mathcal{H}u=0$ on a neighborhood of $Q_{2R}(t_{0},x_{0})$, respectively $\mathcal{H}^{\star}v=0$ and $Q_{2R}^{\star}(t_{0},x_{0})$, respectively, have local bounds of the form, respectively,
@@
Up to changing the constant $B$ and the scale, the two notions are equivalent. More precisely, if $\mathcal{H}$ satisfies Moser’s $L^{2}$-$L^{\infty}$ estimates then it satisfies the local boundedness property. Conversely, if $\mathcal{H}$ satisfies the local boundedness property then it satisfies Moser’s $L^{2}$-$L^{\infty}$ estimates with a scale of $3R$ instead of $2R$ in (5.2). The same statement applies to $\mathcal{H}^{\star}$. This follows easily from (4.1) and the doubling property (2.3).
@@
The conditions (5.2) and (5.3) are usually presented by taking suprema on $Q_{R}(t_{0},x_{0})$ and $Q_{R}^{\star}(t_{0},x_{0})$ respectively, which means that one would need to know a priori that local weak solutions have pointwise values. In contrast, Moser’s $L^{2}$-$L^{\infty}$ estimates offer a weaker formulation, as they only require taking the essential supremum over $Q_{R}(t_{0},x_{0})$ and $Q_{R}^{\star}(t_{0},x_{0})$. In particular, it does not require knowing that solutions are continuous, or even defined, at every point. This more relaxed formulation, or more precisely the local boundedness property (5.4) and (5.5), which is an equivalent formulation, has been used in kinetic and parabolic contexts in [AIN24] and [AE23], respectively.
@@
If $\mathcal{H}$ and $\mathcal{H}^{\star}$ satisfy Moser’s $L^{2}$-$L^{\infty}$ estimates, then $\mathcal{H}$ has Gaussian upper bounds.
@@
This also applies to the adjoint $\tilde{\Gamma}(s,t)$. In fact, a similar computation shows
$\|e^{\psi}\tilde{\Gamma}(s,t)e^{-\psi}f\|^{2}_{2,\omega}\leq e^{\frac{2M^{2}\gamma^{2}}{\nu}(t-s)}\left\|f\right\|^{2}_{2,\omega}.$
Using Moser’s $L^{2}$-$L^{\infty}$ estimates with $R=\frac{\sqrt{t-s}}{2}$, we have for almost every $x\in\mathbb{R}^{n}$,
$\left|U(t,x)\right|^{2}\leq\frac{B^{2}}{\mu(Q_{\sqrt{t-s}}(t,x))}\int_{s}^{t}\int_{B(x,\sqrt{t-s})}\left|U(\tau,y)\right|^{2}\ \mathrm{d}\mu,$
@@
$|e^{\psi(x)}U(t,x)|^{2}$ $\leq\frac{B^{2}}{(t-s)\omega_{t-s}(x)}\int_{s}^{t}\int_{B(x,\sqrt{t-s})}e^{2(\psi(x)-\psi(y)}|e^{\psi(y)}U(\tau,y)|^{2}\ \mathrm{d}\omega(y)\mathrm{d}\tau$
$\leq\frac{B^{2}e^{2\gamma\sqrt{t-s}}}{(t-s)\omega_{t-s}(x)}\int_{s}^{t}\int_{B(x,\sqrt{t-s})}|e^{\psi(y)}U(\tau,y)|^{2}\ \mathrm{d}\omega(y)\mathrm{d}\tau$
$\leq\frac{B^{2}e^{2\gamma\sqrt{t-s}}}{(t-s)\omega_{t-s}(x)}\int_{s}^{t}\|e^{\psi}U(\tau)\|^{2}_{2,\omega}\ \mathrm{d}\tau.$
Using (5.6), we have for almost every $x\in\mathbb{R}^{n}$,
(5.7) $|e^{\psi(x)}U(t,x)|^{2}\leq\frac{B^{2}e^{2\gamma\sqrt{t-s}}}{(t-s)\omega_{t-s}(x)}\left(\int_{s}^{t}e^{\frac{2M^{2}\gamma^{2}}{\nu}(\tau-s)}\ \mathrm{d}\tau\right)\left\|f\right\|^{2}_{2,\omega}.$
If $\gamma=0$, we have
$\left|U(t,x)\right|^{2}\leq\frac{B^{2}}{\omega_{t-s}(x)}\left\|f\right\|^{2}_{2,\omega}.$
Hence,
(5.8) $\left\|\sqrt{\omega_{t-s}}\ \Gamma(t,s)f\right\|_{L^{\infty}(\mathbb{R}^{n})}\leq B\left\|f\right\|_{2,\omega}.$
@@
(5.9) $\|\sqrt{\omega_{t-s}}\ \tilde{\Gamma}(s,t)f\|_{L^{\infty}(\mathbb{R}^{n})}\leq B\left\|f\right\|_{2,\omega}.$
@@
$|e^{\psi(x)}U(t,x)|^{2}\leq\frac{B^{2}e^{2\gamma\sqrt{t-s}}}{(t-s)\omega_{t-s}(x)}\frac{e^{2M^{2}\gamma^{2}(t-s)/\nu}}{2M^{2}\gamma^{2}/\nu}\left\|f\right\|^{2}_{2,\omega}.$
@@
$\|\sqrt{\omega_{t-s}}\ e^{\psi}\Gamma(t,s)e^{-\psi}f\|_{L^{\infty}(\mathbb{R}^{n})}\leq\frac{B}{M\sqrt{2/\nu}}\frac{e^{\gamma\sqrt{t-s}+\frac{M^{2}\gamma^{2}}{\nu}(t-s)}}{\gamma\sqrt{t-s}}\left\|f\right\|_{2,\omega}.$
@@
(5.10) $\|\sqrt{\omega_{t-s}}\ e^{\psi}\tilde{\Gamma}(s,t)e^{-\psi}f\|_{L^{\infty}(\mathbb{R}^{n})}\leq\frac{B}{M\sqrt{2/\nu}}\frac{e^{\gamma\sqrt{t-s}+\frac{M^{2}\gamma^{2}}{\nu}(t-s)}}{\gamma\sqrt{t-s}}\left\|f\right\|_{2,\omega}.$
@@
(5.11) $\|\Gamma(t,s)(\sqrt{\omega_{t-s}}f)\|_{2,\omega}\leq B\left\|f\right\|_{1,\omega}.$
@@
(5.12) $\|e^{\psi}\Gamma(t,s)(\sqrt{\omega_{t-s}}\ e^{-\psi}f)\|_{2,\omega}\leq\frac{B}{M\sqrt{2/\nu}}\frac{e^{\gamma\sqrt{t-s}+\frac{M^{2}\gamma^{2}}{\nu}(t-s)}}{\gamma\sqrt{t-s}}\left\|f\right\|_{1,\omega},$
@@
$\sqrt{\omega_{t-s}}$ $e^{\psi}\Gamma(t,s)(\sqrt{\omega_{t-s}}\ e^{-\psi}f)=\sqrt{\omega_{t-s}}\ e^{\psi}\Gamma(t,\frac{t+s}{2})e^{-\psi}(e^{\psi}\Gamma(\frac{t+s}{2},t)\sqrt{\omega_{t-s}}\ e^{-\psi}f)$
@@
(5.14) $\left\|\sqrt{\omega_{t-s}}\ \Gamma(t,s)(\sqrt{\omega_{t-s}}f)\right\|_{L^{\infty}(\mathbb{R}^{n})}\leq B^{2}D\left\|f\right\|_{1,\omega}.$
@@
(5.15) $\|\sqrt{\omega_{t-s}}\ e^{\psi}\Gamma(t,s)(\sqrt{\omega_{t-s}}\ e^{-\psi}f)\|_{L^{\infty}(\mathbb{R}^{n})}\leq\frac{B^{2}D\nu}{M^{2}}\frac{e^{\gamma\sqrt{2(t-s)}+\frac{M^{2}\gamma^{2}}{\nu}(t-s)}}{\gamma^{2}(t-s)}\left\|f\right\|_{1,\omega},$
@@
(5.16) $\left|\Gamma(t,x;s,y)\right|\leq\frac{1}{\sqrt{\omega_{t-s}(x)}\sqrt{\omega_{t-s}(y)}}B^{2}D.$
@@
(5.17) $\left|\Gamma(t,x;s,y)\right|\leq\frac{B^{2}D\nu/M^{2}}{\sqrt{\omega_{t-s}(x)}\sqrt{\omega_{t-s}(y)}}\times\frac{\exp(\gamma\sqrt{2(t-s)}+\frac{M^{2}\gamma^{2}}{\nu}(t-s))}{\gamma^{2}(t-s)}\times\exp(\psi(y)-\psi(x)).$
@@
(5.18) $\left|\Gamma(t,x;s,y)\right|\leq\frac{B^{2}D\nu/M^{2}}{\sqrt{\omega_{t-s}(x)}\sqrt{\omega_{t-s}(y)}}\times\frac{\exp(\gamma\sqrt{2(t-s)}+\kappa\gamma^{2}(t-s)-\gamma\left|x-y\right|)}{\gamma^{2}(t-s)}.$
@@
$\frac{\exp(\gamma\sqrt{2(t-s)}+\kappa\gamma^{2}(t-s)-\gamma\left|x-y\right|)}{\gamma^{2}(t-s)}=4\kappa^{2}\frac{\exp(\frac{\xi}{\sqrt{2}\kappa}-\frac{\xi^{2}}{4\kappa})}{\xi^{2}}.$
@@
(5.19) $\left|\Gamma(t,x;s,y)\right|\leq\frac{B^{2}D}{\sqrt{\omega_{t-s}(x)}\sqrt{\omega_{t-s}(y)}}\min\left(1,4\frac{M^{2}}{\nu}\times\frac{\exp(\frac{\xi}{\sqrt{2}\kappa}-\frac{\xi^{2}}{4\kappa})}{\xi^{2}}\right).$
@@
$4\frac{M^{2}}{\nu}\cdot\frac{\exp(\frac{\xi^{\prime}}{\sqrt{2}\kappa}-\frac{{\xi}^{\prime 2}}{4\kappa})}{{\xi}^{\prime 2}}\leq\exp(-\frac{{\xi^{\prime}}^{2}}{8\kappa})\cdot$
@@
$\left|\Gamma(t,x;s,y)\right|\leq\frac{B^{2}D}{\sqrt{\omega_{t-s}(x)}\sqrt{\omega_{t-s}(y)}}\exp(-\frac{\nu}{8M^{2}}\frac{{\left|x-y\right|}^{2}}{(t-s)}).$
@@
$\left|\Gamma(t,x;s,y)\right|\leq\frac{B^{2}D}{\sqrt{\omega_{t-s}(x)}\sqrt{\omega_{t-s}(y)}}e^{R^{2}}\exp\left(-\frac{\left|x-y\right|^{2}}{t-s}\right).$
@@
If $\mathcal{H}$ has Gaussian upper bounds, then $\mathcal{H}$ and $\mathcal{H}^{\star}$ satisfy Moser’s $L^{2}$-$L^{\infty}$ estimates.
@@
Estimate of $\left\|v_{2}(t,\cdot)\right\|_{L^{\infty}(B(x_{0},R))}$: likewise, since $-\omega^{-1}A(t,\cdot)\nabla_{x}u\cdot\nabla_{x}\zeta$ is compactly supported in time and space, we have by Cauchy-Schwarz inequality and for almost every $x\in B(x_{0},R)$,
@@
[…]

### v3

@@ 4.3. Conservation property
When $R\to\infty$, the term in the left hand side tends to $\int_{\mathbb{R}^{n}}\overline{\psi(x)}\ \mathrm{d}\omega(x)$ and the first term in the right hand side tends to $\int_{\mathbb{R}^{n}}\overline{(\tilde{\Gamma}(s,t)\psi)(x)}\ \mathrm{d}\omega(x)$ by Corollary 4.2. It remains to show that the second term in the right hand side tends to 0 as $R\to\infty$. For $R>1$, we have
@@
5. Gaussian upper bounds and Moser’s $^{2}-LL^{\infty}$ estimates
In this section, we prove the equivalence between Gaussian upper bounds for $\mathcal{H}$ and Moser’s $^{2}-LL^{\infty}$ estimates for $\mathcal{H}$ and $\mathcal{H}^{\star}$. We will first define these notions rigorously.
@@ 5.1. Gaussian upper bounds and the generalized fundamental solution
See [CR14, Rem. 3], where it is proven that there is a constant $C=C(D,k_{0})>0$ verifying
@@
5.2. Moser’s $^{2}-LL^{\infty}$ estimates
@@
We say that $\mathcal{H}$, respectively $\mathcal{H}^{\star}$, satisfies Moser’s $^{2}-LL^{\infty}$ estimates if there exists a constant $B>0$ such that that for all $R>0$, $(t_{0},x_{0})\in\mathbb{R}^{1+n}$, and all local weak solutions of $\mathcal{H}u=0$ on a neighborhood of $Q_{2R}(t_{0},x_{0})$, respectively $\mathcal{H}^{\star}v=0$ and $Q_{2R}^{\star}(t_{0},x_{0})$, respectively, have local bounds of the form, respectively,
@@
Up to changing the constant $B$ and the scale, the two notions are equivalent. More precisely, if $\mathcal{H}$ satisfies Moser’s $^{2}-LL^{\infty}$ estimates then it satisfies the local boundedness property. Conversely, if $\mathcal{H}$ satisfies the local boundedness property then it satisfies Moser’s $^{2}-LL^{\infty}$ estimates with a scale of $3R$ instead of $2R$ in (5.6). The same statement applies to $\mathcal{H}^{\star}$. This follows easily from (4.3) and the doubling property (2.3).
@@
The conditions (5.6) and (5.7) are usually presented by taking suprema on $Q_{R}(t_{0},x_{0})$ and $Q_{R}^{\star}(t_{0},x_{0})$ respectively, which means that one would need to know a priori that local weak solutions have pointwise values. In contrast, Moser’s $^{2}-LL^{\infty}$ estimates offer a weaker formulation, as they only require taking the essential supremum over $Q_{R}(t_{0},x_{0})$ and $Q_{R}^{\star}(t_{0},x_{0})$. In particular, it does not require knowing that solutions are continuous, or even defined, at every point. This more relaxed formulation, or more precisely the local boundedness property (5.9) and (5.10), which is an equivalent formulation, has been used in kinetic and parabolic contexts in [AIN25] and [AE23], respectively.
@@
If $\mathcal{H}$ and $\mathcal{H}^{\star}$ satisfy Moser’s $^{2}-LL^{\infty}$ estimates, then $\mathcal{H}$ has Gaussian upper bounds.
@@
This also applies to the adjoint $\tilde{\Gamma}(s,t)$. In fact, a similar computation (or duality) shows
$\|e^{-\psi}\tilde{\Gamma}(s,t)e^{\psi}f\|^{2}_{2,\omega}\leq e^{\frac{2M^{2}\gamma^{2}}{\nu}(t-s)}\left\|f\right\|^{2}_{2,\omega}.$
Moser’s $^{2}LL^{\infty}$ estimate (5.6) with $R=\frac{\sqrt{t-s}}{2}$, together with (4.3), implies that for all $x\in\mathbb{R}^{n}$ and for almost every $z\in B(x,R)$,
$\left|U(t,z)\right|^{2}\leq\frac{B^{2}}{\mu(Q_{\sqrt{t-s}}(t,x))}\int_{s}^{t}\int_{B(x,\sqrt{t-s})}\left|U(\tau,y)\right|^{2}\ \mathrm{d}\mu.$
Hence, for almost every $z\in B(x,R)$,
$|e^{\psi(z)}U(t,z)|^{2}$ $\leq\frac{B^{2}}{(t-s)\omega_{t-s}(x)}\int_{s}^{t}\int_{B(x,\sqrt{t-s})}e^{2(\psi(z)-\psi(y)}|e^{\psi(y)}U(\tau,y)|^{2}\ \mathrm{d}\omega(y)\mathrm{d}\tau$
$\leq\frac{B^{2}e^{4\gamma\sqrt{t-s}}}{(t-s)\omega_{t-s}(x)}\int_{s}^{t}\int_{B(x,\sqrt{t-s})}|e^{\psi(y)}U(\tau,y)|^{2}\ \mathrm{d}\omega(y)\mathrm{d}\tau$
$\leq\frac{B^{2}e^{4\gamma\sqrt{t-s}}}{(t-s)\omega_{t-s}(x)}\int_{s}^{t}\|e^{\psi}U(\tau)\|^{2}_{2,\omega}\ \mathrm{d}\tau.$
Remark that for all $z\in B(x,R)$, we have $B(z,\frac{\sqrt{t-s}}{2})\subset B(x,\sqrt{t-s})$, and by using the doubling property (2.3), we obtain
$\omega_{t-s}(z)=\omega(B(z,\sqrt{t-s}))\leq D\,\omega(B(z,\frac{\sqrt{t-s}}{2}))\leq D\omega_{t-s}(x).$
Thus, for almost every $z\in B(x,R)$,
$|e^{\psi(z)}U(t,z)|^{2}\leq\frac{DB^{2}e^{4\gamma\sqrt{t-s}}}{(t-s)\omega_{t-s}(z)}\int_{s}^{t}\|e^{\psi}U(\tau)\|^{2}_{2,\omega}\ \mathrm{d}\tau.$
Using this together with (5.13), we have for almost every $x\in\mathbb{R}^{n}$,
(5.14) $|e^{\psi(x)}U(t,x)|^{2}\leq\frac{DB^{2}e^{4\gamma\sqrt{t-s}}}{(t-s)\omega_{t-s}(x)}\left(\int_{s}^{t}e^{\frac{2M^{2}\gamma^{2}}{\nu}(\tau-s)}\ \mathrm{d}\tau\right)\left\|f\right\|^{2}_{2,\omega}.$
If $\gamma=0$, we have
$\left|U(t,x)\right|^{2}\leq\frac{DB^{2}}{\omega_{t-s}(x)}\left\|f\right\|^{2}_{2,\omega}.$
@@
(5.15) $\left\|\sqrt{\omega_{t-s}}\ \Gamma(t,s)f\right\|_{L^{\infty}(\mathbb{R}^{n})}\leq D^{1/2}B\left\|f\right\|_{2,\omega}.$
@@
(5.16) $\|\sqrt{\omega_{t-s}}\ \tilde{\Gamma}(s,t)f\|_{L^{\infty}(\mathbb{R}^{n})}\leq D^{1/2}B\left\|f\right\|_{2,\omega}.$
@@
$|e^{\psi(x)}U(t,x)|^{2}\leq\frac{DB^{2}e^{4\gamma\sqrt{t-s}}}{(t-s)\omega_{t-s}(x)}\frac{e^{2M^{2}\gamma^{2}(t-s)/\nu}}{2M^{2}\gamma^{2}/\nu}\left\|f\right\|^{2}_{2,\omega}.$
@@
$\|\sqrt{\omega_{t-s}}\ e^{\psi}\Gamma(t,s)e^{-\psi}f\|_{L^{\infty}(\mathbb{R}^{n})}\leq\frac{D^{1/2}B}{M\sqrt{2/\nu}}\frac{e^{2\gamma\sqrt{t-s}+\frac{M^{2}\gamma^{2}}{\nu}(t-s)}}{\gamma\sqrt{t-s}}\left\|f\right\|_{2,\omega}.$
@@
(5.17) $\|\sqrt{\omega_{t-s}}\ e^{-\psi}\tilde{\Gamma}(s,t)e^{\psi}f\|_{L^{\infty}(\mathbb{R}^{n})}\leq\frac{D^{1/2}B}{M\sqrt{2/\nu}}\frac{e^{2\gamma\sqrt{t-s}+\frac{M^{2}\gamma^{2}}{\nu}(t-s)}}{\gamma\sqrt{t-s}}\left\|f\right\|_{2,\omega}.$
@@
(5.18) $\|\Gamma(t,s)(\sqrt{\omega_{t-s}}f)\|_{2,\omega}\leq D^{1/2}B\left\|f\right\|_{1,\omega},$
@@
(5.19) $\|e^{\psi}\Gamma(t,s)(\sqrt{\omega_{t-s}}\ e^{-\psi}f)\|_{2,\omega}\leq\frac{D^{1/2}B}{M\sqrt{2/\nu}}\frac{e^{2\gamma\sqrt{t-s}+\frac{M^{2}\gamma^{2}}{\nu}(t-s)}}{\gamma\sqrt{t-s}}\left\|f\right\|_{1,\omega},$
@@
$\sqrt{\omega_{t-s}}\$$e^{\psi}\Gamma(t,s)(\sqrt{\omega_{t-s}}\ e^{-\psi}f)=\sqrt{\omega_{t-s}}\ e^{\psi}\Gamma(t,\frac{t+s}{2})e^{-\psi}(e^{\psi}\Gamma(\frac{t+s}{2},t)\sqrt{\omega_{t-s}}\ e^{-\psi}f)$
@@
(5.21) $\left\|\sqrt{\omega_{t-s}}\ \Gamma(t,s)(\sqrt{\omega_{t-s}}f)\right\|_{L^{\infty}(\mathbb{R}^{n})}\leq B^{2}D^{2}\left\|f\right\|_{1,\omega}.$
@@
(5.22) $\|\sqrt{\omega_{t-s}}\ e^{\psi}\Gamma(t,s)(\sqrt{\omega_{t-s}}\ e^{-\psi}f)\|_{L^{\infty}(\mathbb{R}^{n})}\leq\frac{B^{2}D^{2}\nu}{M^{2}}\frac{e^{2\gamma\sqrt{2(t-s)}+\frac{M^{2}\gamma^{2}}{\nu}(t-s)}}{\gamma^{2}(t-s)}\left\|f\right\|_{1,\omega},$
@@
(5.23) $\left|\Gamma(t,x;s,y)\right|\leq\frac{1}{\sqrt{\omega_{t-s}(x)}\sqrt{\omega_{t-s}(y)}}B^{2}D^{2}.$
@@
(5.24) $\left|\Gamma(t,x;s,y)\right|\leq\frac{B^{2}D^{2}\nu/M^{2}}{\sqrt{\omega_{t-s}(x)}\sqrt{\omega_{t-s}(y)}}\times\frac{\exp(2\gamma\sqrt{2(t-s)}+\frac{M^{2}\gamma^{2}}{\nu}(t-s))}{\gamma^{2}(t-s)}\times\exp(\psi(y)-\psi(x)).$
@@
(5.25) $\left|\Gamma(t,x;s,y)\right|\leq\frac{B^{2}D^{2}\nu/M^{2}}{\sqrt{\omega_{t-s}(x)}\sqrt{\omega_{t-s}(y)}}\times\frac{\exp(2\gamma\sqrt{2(t-s)}+\kappa\gamma^{2}(t-s)-\gamma\left|x-y\right|)}{\gamma^{2}(t-s)}.$
@@
$\frac{\exp(2\gamma\sqrt{2(t-s)}+\kappa\gamma^{2}(t-s)-\gamma\left|x-y\right|)}{\gamma^{2}(t-s)}=4\kappa^{2}\frac{\exp(\frac{\sqrt{2}}{\kappa}\xi-\frac{1}{4\kappa}\xi^{2})}{\xi^{2}}.$
@@
(5.26) $\left|\Gamma(t,x;s,y)\right|\leq\frac{B^{2}D^{2}}{\sqrt{\omega_{t-s}(x)}\sqrt{\omega_{t-s}(y)}}\min\left(1,\frac{4M^{2}}{\nu}\frac{\exp(\frac{\sqrt{2}}{\kappa}\xi-\frac{1}{4\kappa}\xi^{2})}{\xi^{2}}\right).$
@@
$\frac{4M^{2}}{\nu}\cdot\frac{\exp(\frac{\sqrt{2}}{\kappa}\xi-\frac{1}{4\kappa}\xi^{2})}{\xi^{2}}\leq\exp(-\frac{{\xi^{\prime}}^{2}}{8\kappa})\cdot$
@@
$\left|\Gamma(t,x;s,y)\right|\leq\frac{B^{2}D^{2}}{\sqrt{\omega_{t-s}(x)}\sqrt{\omega_{t-s}(y)}}\exp(-\frac{\nu}{8M^{2}}\frac{{\left|x-y\right|}^{2}}{(t-s)}).$
@@
$\left|\Gamma(t,x;s,y)\right|\leq\frac{B^{2}D^{2}}{\sqrt{\omega_{t-s}(x)}\sqrt{\omega_{t-s}(y)}}e^{R^{2}}\exp\left(-\frac{\left|x-y\right|^{2}}{t-s}\right).$
@@
If $\mathcal{H}$ has Gaussian upper bounds, then $\mathcal{H}$ and $\mathcal{H}^{\star}$ satisfy Moser’s $^{2}-LL^{\infty}$ estimates.
@@
Estimate of $\left\|v_{2}(t,\cdot)\right\|_{L^{\infty}(B(x_{0},R))}$: likewise, since $\omega^{-1}A(t,\cdot)\nabla_{x}u\cdot\nabla_{x}\zeta$ is compactly supported in time and space, we have by Cauchy-Schwarz inequality and for almost every $x\in B(x_{0},R)$,
@@
$\int_{\Sigma^{\prime}}|\tilde{\Gamma}(s,t)\phi(y)|^{2}\ \mathrm{d}\mu(s,y)$ $=\int_{\Sigma^{\prime}}\left|\int_{B(x_{0},R)}\overline{\Gamma}(t,z;s,y)\phi(z)\ \mathrm{d}\omega(z)\right|^{2}d\mu(s,y)$